Set up and solve a system of linear equations for the the following:
find a, b and c for the quadratic function: f(x)=ax^2+bx+c
Given that
f(x)=1, f(-3)=17, f(2)=-1/2
I have completed the quadratic function portion but I'm really lost finding a, b and c...Pleese help.

Answer by longjonsilver(2297)

f(-3)=17 means when x=-3, y=17
f(2)=-1/2 means when x=2, y=-1/2.

Put these 2 sets of information into the function to get 2 equations. However, as there are 3unknowns (a, b and c) you will need a third equation too: f(x)=1?? is this a typo on your part? shouldn't the x be a number?

Once you have the 3 equations, you will have to solve them simultaneously either by substitution or elimination.

Jon.